The time it takes to complete this assignment varies Normally with mean 20 minutes and standard deviation 2.6 minutes. What is the probability that it takes you at least 21 minutes to complete this assignment?

CAREFUL: YOU NEED TO USE THE TABLE TO GET THIS QUESTION CORRECT. Round your z-score to 2 decimal places before using the table.
Answer #1

Solution :

Given that,

mean = = 20

standard deviation = =2.6

P(x > 21) = 1 - P(x<21 )

= 1 - P[(x -) / < (21-20) /6 ]

= 1 - P(z <0.17 )

Using z table

= 1 -  0.5675

probability= 0.4325